HI i'm a little stuck on this current problem of mine. Would love some help and insight. thank you in adavance :)

Suppose that a household spent $350 less in 2015 by purchasing its cars and home
insurances from Company A than it had done in 2014 by purchasing the cars
insurances from Company B and the home insurance from Company C. Given that
 the household had spent a maximum of $2,000 on insurances in 2015, and
 the household spent $50 less on cars insurances to Company B than on
home insurance to Company C in 2014,
write an inequality to represent the amount that the total household spent to
Companies B and C in 2014. Hence, using the inequalities found, determine the
maximum amount the household spent to each of these companies in 2014

Answer by stanbon(75887)

Suppose that a household spent $350 less in 2015 by purchasing its cars and home
insurances from Company A than it had done in 2014 by purchasing the cars
insurances from Company B and the home insurance from Company C. Given that
 the household had spent a maximum of $2,000 on insurances in 2015, and
 the household spent $50 less on cars insurances to Company B than on
home insurance to Company C in 2014,
------
Write the above using symbols::
A = B+C-350
A <= 2000
B = C-50
-------------------
write an inequality to represent the amount that the total household spent to
Companies B and C in 2014.
B+C -350 <= 2000
B+C <= 2350
---------------
Hence, using the inequalities found, determine the
maximum amount the household spent to each of these companies in 2014
Substitute for "B" to get:
C-50 + C <= 2350
2C <= 2400
C <= $1200
---
Then B <= 2350-1200 = $1150
==========
Cheers,
Stan H.
----------